OpenStudy (anonymous):

HELP!!! Find the area of the portion of the circle that is outside triangle RST if the radius of the circle is 8 cm

OpenStudy (anonymous):

heres the diagram

OpenStudy (gudden):

@denizen .. just think a little over it... to find.. the angles.. in the center... first....

OpenStudy (anonymous):

the angles are 60, 70 and 50 right

OpenStudy (gudden):

first ol all... tell me the way u want to solve it...

OpenStudy (gudden):

cuz.. it appears.. that we are thinking over it in different ways..

OpenStudy (gudden):

@surjithayer .. what are u typing bro...? :p

OpenStudy (anonymous):

i am thinking getting the area of the circle minus the area of the triangle

OpenStudy (gudden):

@denizen take a look on the following links im sure that they are gonna help u alot in solving the problem http://www.mathsisfun.com/geometry/circle-sector-segment.html http://mathschallenge.net/library/geometry/circle_properties

OpenStudy (gudden):

enjoy the question...!.. just give it a little time.. n am sure that u would get the answer.. n its... not that damn tough as it looks like..

OpenStudy (anonymous):

\[\cos 120=\frac{ 8^{2}+8^{2}-(ST)^{2} }{ 2*8*8 }\] \[\cos \left( 180-60 \right)*128=128-(ST)^{2}\] \[(-\cos 60)*128-128=-(ST)^{2}\] \[-\frac{ 1 }{ 2 }*128-128=-(ST)^{2},(ST)^{2}=192=3*64\] \[ST=8\sqrt{3}\] Use sine formula to find RT \[\frac{ RT }{ \sin 70 }=\frac{ ST }{ \sin 60 }\] \[RT=\frac{ ST }{ \frac{ \sqrt{3} }{ 2 } }\sin 70=8\sqrt{3}*\frac{ 2 }{ \sqrt{3} }\sin 70=16 \sin 70\]=15.04 \[\frac{ RS }{ \sin 50 }=\frac{ ST }{ \sin 60 }=\frac{ 8\sqrt{3} }{ \frac{ \sqrt{3} }{ 2 } }=16\] Rs=16 sin50=12.26 \[\frac{ SP }{ RS }=\sin 60=\frac{ \sqrt{3} }{ 2 },SP=\frac{ \sqrt{3} }{ 2 }*RS=\frac{ \sqrt{3} }{ 2 }*12.26=\sqrt{3}*6.13\] \[\sqrt{3}*6.13=10.62\] SP=10.62 Area of triangle\[=\frac{ 1 }{ 2 }*RT*SP=\frac{ 1 }{ 2 }*15.04*10.62=79.86\] \[area of circle=\pi r ^{2}=\pi 8^{2}=64 \pi \] required area\[=64\pi-79.86=201.06-79.86=121.2 sq. cm\]

OpenStudy (anonymous):